Show HOW to use inverses to algebraically solve for x: e2x + 5e* - 14 = 0
Answer #1

p2x sea_14 =0 54-14 EO = y + 5y-1420 y² + 7y-ay-1450 y (4+7) - 2(3+2) == (y +7 (y-2) = 0 = y +7=0, 4 = 2 YE-7 ex=-7 = 2 ? -2